If you have 6 quarts of a 60% acid solution, how many quarts of water would you add to obtain a 40% acid solution?
Answer by stanbon(75887)

If you have 6 quarts of a 60% acid solution, how many quarts of water would you add to obtain a 40% acid solution?
-----------------------------------------
Equation:
acid + acid = acid
0.60*6 + 0*x = 0.40(6+x)
Multiply thru by 100 to get:
60*6 + 0 = 40*6 + 40x
40x = 20*6
x = 3 quarts (amt. of water needed for the mixture)
================
Cheers,
Stan H.